Grade 12Mechanics

A car is moving towards south with a speed of 20 m/s. A motorcyclist is moving towards east with speed of 15 ms–1. At certain instant, the motorcyclist is due south of the car and is at a distance of 50 m from the car. Then time after which they are closest is —(Answer Choices are a) 1/3 s b) 8/3 s c) 1/5 s d) 8/5 s )

Aishwarya
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
 

Dear student, 

let t be the time from the instant when the bike is due south. distance squared = (50-20t)2 +(15t)2= 625t2-2000t +2500

='s'. ds/dt=0, 1250t-2000=0, tmin=1.6 seconds, d2s/dt2=1250t=1250*1.6>0, minimum distance=√s at t=1.6 =√900=30m